# Multiphoton Spectroscopy of a Dynamical Axion Insulator

### v2

@@ Abstract
I Introduction
Axionic particles were originally proposed in high-energy physics in order to solve the charge-parity problem as they were associated to the Nambu-Goldstone bosons of a new global axial U(1) symmetry [1, 2]. Today, axions are a leading dark matter candidate and could help explain the matter-antimatter asymmetry present in the universe [3]. While the axion has yet to be discovered in particle physics, its condensed matter counterpart has been theorized to exist in certain three-dimensional materials [4, 5, 6, 7, 8, 9, 10]. A characteristic identifier of the presence of an axion is a signature magnetoelectric response, leading to a modified form of Maxwell’s equations known as axion electrodynamics [11, 12, 13]; these give rise to physical phenomena which can broadly be classified as either transport or density responses. Axion-induced transport includes anomalous Hall currents which are related to the gradient of the axion, and a magnetic conductivity that relates to the axion’s first time derivative. Density responses have to do with the topological magnetoelectric effect where the axion acts as the coupling constant between magnetic field and polarization; and similarly between electric field and magnetization.
In correlated systems, the axion field has an associated collective excitation which couples nonlinearly to the electromagnetic field; this induces a dynamical magnetoelectric effect on top of the mean field response due to the background axion field [14, 15, 16, 17]. A variety of candidate materials have been put forward where axion field effects manifest in the electronic band structure. For example, axion fields naturally appear in correlated Weyl semimetals that undergo lattice translation symmetry breaking due to the emergence of charge density wave (CDW) order. In this case, the mean background axion field is determined by the separation of Weyl points in momentum space and in energy; while its dynamical contribution is associated with collective excitations of the phase of the CDW, i.e., the Nambu-Goldstone bosons [18, 19]. In the presence of uncompensated carrier densities the softening of the axionic collective mode may even lead to spatially inhomogeneous correlated phases [20]. 3D topological insulators (TI) can also support a static, background axion field which, due to time-reversal symmetry, acts as a $\mathbb{Z}_{2}$ topological invariant and takes quantized values $\theta=0$ or $\pi$. When interactions are turned on that spontaneously break time-reversal symmetry, as in the case of antiferromagnetic (AFM) order, $\theta$ is promoted to a dynamical quantity that takes continuous values between 00 and $2\pi$ [5]. Here, the fluctuations of the axion field are induced by the spin-wave excitations of the AFM order parameter [5, 21, 22].
@@ I Introduction
Here, we define an all-optical, contactless protocol for the identification of nonlinear signatures of dynamical axion excitations in a three-dimensional insulating system, termed dynamical axion insulator (DAI). We keep the approach general such that the results can be applied to any system which hosts axionic excitations. This protocol breaks down into two parts; first a dynamical axion mode is resonantly excited from a background axion field through two-photon Raman processes. Specifically, by solving the dynamics of electromagnetic fields coupled nonlinearly to collective axion modes we show that the transient evolution of the latter carries a distinct dependence on the incident angle, polarization, and frequencies of the two photons. Following excitation, the induced axion electrodynamics are then detected through a third probe beam, which monitors the oscillations of the Kerr angle and acts as a faithful proxy for tracking the dynamical evolution of the axion field in the material [26, 27]. We apply our protocol to a minimal Hamiltonian model that serves as the basis of describing a variety of candidate materials, e.g., Weyl semimetals with CDW order such as (TaSe4)2[23], or magnetic axion insulators such as Co3Sn2S2 [28] and BiFe2Se3 [5]. The study of axionic collective modes in condensed matter systems offers both fundamental insights into their high-energy analogues, as well as practical applications, e.g., in nonreciprocal thermal emitters or rectifiers [29, 30]. In addition, our work demonstrates the potential of exploiting entanglement pairs to probe correlated phases and engineer strong nonlinear optical responses in quantum materials.
Axion electrodynamics
We start by considering the axion-modified electromagnetic Lagrangian
@@
where $\kappa$ is a materials-dependent parameter, to be given a physical interpretation as “compressibility” later on when we consider the specific case of an axion insulator. We have also included a finite gap for the collective mode $\sim\Omega_{0}$, which reflects the possibility of a resonance at finite frequency, e.g., due to collective mode pinning from disorder. As we are dealing with electrodynamics, we assume that the speed at which the axion propagates is orders of magnitude slower than the speed of light; as such the axion is effectively dispersionless provided it has a finite gap $\Omega_{0}$. Only in the case of the gap $\Omega_{0}$ going to zero would the axion dispersion become important. This can easily be seen from the axion dispersion, given as $\omega^{2}=(\nu k)^{2}+\Omega_{0}^{2}$ for frequency $\omega$, speed $\nu$, and momentum $k$ of the axion. Comparing to the speed of light, $c\gg\nu$ so that all terms relating to first gradients of the dynamical axion field, $\nabla\delta\theta$, can be safely ignored; these describe effects from axion dispersion through the medium.
@@
where we have included a phenomenologically relevant dissipation term $\kappa\gamma\partial_{t}\delta\theta$, with damping rate $\gamma$, which represents the coupling of the order parameter to other degrees of freedom in the bulk. The above equation describes the dynamics of a massive spinless bosonic field, driven by an effective force $\frac{\alpha}{\pi}\bm{E}\cdot\bm{B}$. Thus as long as the electric and magnetic fields have some non-orthogonal component the dynamical axion mode can be excited. In the following, we show how the solutions of these nonlinear coupled equations depend on the different configurations of the driving term $\bm{E}\cdot\bm{B}$ and how can this be utilized in selectively exciting the dynamical axion.
II Results
II.1 Dynamical axion generation
We now describe the protocol to excite a dynamical axion mode, where the approach is kept general and may apply to any axionic, correlated system which may endow the axion field with intrinsic dynamics. For ease of calculation we switch to the electromagnetic gauge field notation which relates the electric and magnetic fields $\bm{E}=-\nabla\phi-\partial_{t}\bm{A}$ and $\bm{B}=\nabla\times\bm{A}$ to the scalar potential $\phi$ and vector potential $\bm{A}$. For simplicity, we take the Weyl gauge where $\phi=0$. Perturbatively expanding $\delta\theta$ and $\bm{A}$ in the small parameter $\alpha$, i.e., the fine structure constant,
$\bm{A}(\bm{x},t)=\bm{A}^{(0)}+\alpha\bm{A}^{(1)}+\alpha^{2}\bm{A}^{(2)}+\alpha^{3}\bm{A}^{(3)}+...,$ (5)
$\delta\theta(\bm{x},t)=\delta\theta^{(0)}+\alpha\delta\theta^{(1)}+\alpha^{2}\delta\theta^{(2)}+\alpha^{3}\delta\theta^{(3)}+...,$ (6)
plugging these expansions into Eq. (4) and collecting terms via their order in $\alpha$, the lowest surviving order to excite a dynamical axion is $\delta\theta^{(1)}$
$\kappa(\partial_{t}^{2}+\gamma\partial_{t}+\Omega_{0}^{2})\delta\theta^{(1)}=\frac{1}{\pi}\partial_{t}\bm{A}^{(0)}(\nabla\times\bm{A}^{(0)}).$ (7)
Performing a Fourier transform of the above expression and simplifying (see Methods section IV.1 for detailed derivation)
$\delta\theta^{(1)}(q)=\frac{1}{\pi\kappa}\left[-\Omega^{2}+i\gamma\Omega+\Omega_{0}^{2}\right]^{-1}\times$
$\int_{q_{1},q_{2}}\delta_{q_{1}+q_{2},q}\omega_{1}\bm{A}^{(0)}(q_{1})\cdot(\bm{q_{2}}\times\bm{A}^{(0)}(q_{2})).$ (8)
Here $\delta\theta^{(1)}(q)$ is the first order dynamical axion response, where $q=(\Omega,\bm{q})$ is the four-momenta, with frequency $\Omega$ and momentum $\bm{q}$. The zeroth order gauge field $\bm{A}^{(0)}(q_{i})$ for $i=1,2$ describes an electromagnetic field with four-momentum $q_{i}=(\omega_{i},\bm{q}_{i})$ for frequency $\omega_{i}$ and momentum $\bm{q}_{i}$.
@@ II.1 Dynamical axion generation
Importantly, the amplitude of oscillations given in Eq. (9) depend on: (i) a geometric volume part defined by the two polarization vectors $\hat{\bm{\varepsilon}}_{i}$ and the three-momentum $q_{i}/\omega_{i}$ of the incident electromagnetic fields, and (ii) a frequency-dependent Lorentzian-type response. As the former is a second order nonlinear optical process, there are essentially two channels for three-wave mixing: two-photon absorption (understood as second harmonic generation or sum frequency generation depending on if the two incident frequencies are equal or not), as shown in a Feynman-like diagram Fig. 2(a); and stimulated Raman scattering (understood as difference frequency generation), as shown in Fig. 2(b). The dependence on the excitation channel can be understood geometrically as the parallelpiped volume made by the polarization unit vectors, $\hat{\varepsilon}_{i}$, and the vector composed of the three-momenta $\bm{q}_{1}/\omega_{1}-\bm{q}_{2}/\omega_{2}$, see Fig. 2(c). Its contribution to the amplitude response of the axion is thus constrained by the allowed geometry of the excitation beams. The excitation scheme that maximizes this volume corresponds to beams with orthogonal polarizations and which are also orthogonal with respect to their relative three-momentum $\bm{q}_{1}/\omega_{1}-\bm{q}_{2}/\omega_{2}$.
We explore the efficacy of this two-photon excitation protocol for a variety of different incident beam angles $\varphi$ and frequencies $\omega_{i}$ in the maximally-crossed polarization geometry, i.e., we keep the polarizations of the excitation beams orthogonal $\hat{\varepsilon}_{1}\cdot\hat{\varepsilon}_{2}=0$, c.f., Fig. 1. In Fig. 3, we show the induced dynamical axion amplitude $|\delta\theta|^{2}$ as a function of incident photon frequencies $\omega_{1},\omega_{2}$ for different incident angles, $\varphi=0,\pi/4$, and $\pi/2$. Depending on the incident angle $\varphi$ made by the beams, we find two distinct mechanisms that can be used to excite the dynamical axion mode. First, when the excitation beams are close to head-on, i.e., $\varphi=0$, the two-photon absorption channel appears in the upper half-plane, as shown in Fig. 3(a); when the excitation frequencies match the axion’s frequency, i.e., $\omega_{1}+\omega_{2}=\Omega_{0}$, the amplitude of oscillations is maximized for this case.
On the other hand, for parallel-propagating collinear beams, i.e., $\varphi=\pi/2$, this channel is entirely suppressed and instead the stimulated (inelastic) Raman excitation channel emerges in the lower half-plane, see Fig. 3(b). Note in this case the relative momentum of the two beams is zero and therefore there must be a finite relative-frequency. As indicated by the color bar, this beam configuration achieves the strongest axion response. This can be understood by again referring to Eq. (9) as well as Fig. 2(c). For configurations where $\omega_{1}>0$ and $\omega_{2}<0$, this will maximize the vector $\bm{q}_{1}/\omega_{1}-\bm{q}_{2}/\omega_{2}$ which will become a sum instead of a difference. In turn this will maximize the parallelpiped volume of Fig. 2(c), as the polarization vectors $\hat{\varepsilon_{i}}$ are kept fixed.
At intermediate incident angles, e.g., $\varphi=\pi/4$, both channels are active, see Fig. 3(c). Due to the dependence of the response to the geometric volume given in Eq. (9), the two-photon absorption channel is significantly weaker as compared to the stimulated Raman scattering channel. We, hence, separate the two half-planes at $y=0$ to resolve these responses. Fig. 3 clearly shows how varying degrees of freedom of the incident beam can excite the axion field, with the inelastic stimulated Raman channel for collinear beams resulting in the strongest response.
@@
Figure 3: This figure shows the two distinct multiphoton absorption pathways with which the axion may be excited, i.e. two-photon absorption in the upper half plane, and stimulated Raman scattering in the lower half plane. Axion excitation amplitude $|\delta\theta|^{2}$ is shown via the color bars. We plot the axion amplitude $|\delta\theta|^{2}$ as a function of the two excitation beam’s frequencies normalized over the axion resonant frequency $\omega_{1}/\Omega_{0}$ and $\omega_{2}/\Omega_{0}$, and their corresponding geometry given at the top of each subfigure. Note the axion resonant frequency is normalized $\Omega_{0}=1$. (a) Excitation amplitude for $\varphi=0$ (beams anti-parallel head-on). (b) Parallel collinear incidence when $\varphi=\pi/2$. In this case, the two-photon absorption channel has been completely suppressed and instead the Raman excitation channel is maximal. (c) Orthogonal incidence when $\varphi=\pi/4$. We see that in addition to the two-photon absorption channel, there is also the emergence of a much stronger inelastic (stimulated Raman) excitation channel. Note the $y=0$ grey line separates the two half planes.
II.2 Axion detection
Now that the protocol for exciting an axion collective mode has been established, we move on to demonstrate how it can be detected via an optical signature. The key idea is to use a linearly polarized detection beam of known frequency at normal incidence, to allow for optical detection of the axion field by its effect on the Kerr angle of the reflected light; a technique well established for the spectroscopy of static axion responses [26].
@@ II.2 Axion detection
The dispersion relation of electromagnetic fields in the material is obtained by solving the equations of motion of the electromagnetic field
$\epsilon\partial^{2}_{t}{\bm{A}}-\frac{1}{\mu}\nabla^{2}\bm{A}=\partial_{t}\bm{J},$ (10)
where $\epsilon$ is the relative permittivity, $\mu$ is the relative permeability, and $\bm{J}$ is the current given in Eq. (2). Without loss of generality we assume a non-magnetic, spatially isotropic medium such that the permeability and permittivity tensors are simplified to $\mu^{-1}=\mathds{1}$ and $\epsilon=\epsilon\mathds{1}$, where $\epsilon$ is a constant characterizing the material. Furthermore, we consider propagation along the z-axis and solve the equations of motion in the plane orthogonal to the direction of propagation. In this geometry, the natural basis is given by $\bm{A}_{\pm}=A_{\pm}\frac{1}{\sqrt{2}}\left(1,\pm i\right)^{T}$ where $A_{\pm}$ is an amplitude prefactor, with corresponding dispersion $k=\omega n_{\pm}$. The refractive index given by (see Methods section IV.2)
$n_{\pm}=\left(\frac{\epsilon\pm\epsilon^{\prime}}{1\pm\mu^{\prime}}\right)^{1/2}$ (11)
where $\epsilon^{\prime}=\omega^{-1}{\partial_{z}\theta}$ is related to the electric conductivity and $\mu^{\prime}=k^{-1}{\partial_{t}\theta}$ is related to the magnetic conductivity.
The reflectance and transmittance amplitudes can be obtained via the continuity of electric and total magnetic fields at the interface, given as $\bm{E}^{I}+\bm{E}^{R}=\bm{E}^{T}$, and $\bm{B}^{I}+\bm{B}^{R}=\bm{B}^{T}+\frac{\alpha}{\pi}\delta\theta\bm{E}^{T}$ (see Methods section IV.2). Without loss of generality, here we only include the dynamical axion $\delta\theta$ contributions to the Fresnel equations, and set the background axion at the boundary to zero. Any such contributions, as in $\mathds{Z}_{2}$ topological insulators where the axion takes quantized values, can be incorporated in the equilibrium position of the dynamical axion mode $\delta\theta$. Using the matching conditions, the coefficients of reflection are computed as
@@
where $R_{\pm}=\frac{A_{\pm}^{R}}{A_{\pm}^{I}}$ is the ratio between the reflected and initial amplitudes of the fields. The reflection coefficients are, in general, different depending on the refractive index of left- and right-circularly polarized light, and the dynamical axion angle. From this, the complex Kerr rotation angle $\Phi_{K}=\Theta_{K}+i\eta_{K}$ may be computed, where the real part $\Theta_{K}$ is termed the polar Kerr rotation, while the imaginary part $\eta_{K}$ is the polar Kerr ellipticity. The polar Kerr rotation relates to the argument of the reflection coefficients, $R_{\pm}=|R_{\pm}|e^{i\Delta_{\pm}}$, as $\Theta_{K}=-\frac{1}{2}(\Delta_{+}-\Delta_{-})$, where $\Delta_{\pm}$ are the phases of the respective reflection coefficient. Meanwhile the polar Kerr ellipticity is related to the amplitudes as $\eta_{K}=(R_{+}-R_{-})/(R_{+}+R_{-})$ [31].
The polar Kerr ellipticity $\eta_{K}$ arises solely from the bulk background axion field and the associated change in the refractive index $n_{\pm}$, see Eq. (11). On the other hand, the polar Kerr rotation $\Theta_{K}$ presented here is entirely distinct from this. It is related to the collective fluctuations of the DAI, therefore, generating a dynamical response of the polar Kerr rotation angle. We note that an additional nontrivial background axion field at the boundary can even lead to a quantized Kerr response [26, 27].
We linearize Eq. (12) for small $\delta\theta$ and focus on the real part of the complex Kerr rotation, $\Theta_{K}$, given by
$\Theta_{K}=-\frac{1}{2}(\Delta_{+}-\Delta_{-})\approx\frac{\alpha}{\pi}\frac{(-2+n_{+}^{2}+n_{-}^{2})}{(n_{+}^{2}-1)(n_{-}^{2}-1)}\delta\theta.$ (13)
The final result of this work is contained within the above equation, which shows the modulation of the reflected beam’s polar Kerr rotation angle and its dependence on the axion field fluctuations, c.f., Eq. (9). This serves as our smoking gun optical signature to unambiguously deduce the presence of a dynamical axion field in the material.
@@
Figure 4: Manifestations of excited axion collective modes in transient Kerr angle rotation. Using the protocol established here, the axion is excited by a pump profile taken as $\bm{E}\cdot\bm{B}\propto e^{(t/\sigma)^{2}}cos(\Omega_{D}t)$, which is tuned to be in two-photon resonance with the dynamical axion, i.e., $\Omega_{D}=\Omega_{0}$. This induces out-of-phase collective oscillations of the axion angle, shown above as $\delta\theta(t)$, with the relative phase determined by the resonant driving condition and the axion dumping parameter $\gamma$; for on-resonance driving it takes the value of $\pi/2$. These oscillations then manifest in the time-dependence of the Kerr-angle $\Theta_{K}(t)$ which approximately track the axion field. We note that the amplitudes depicted here are not to scale.
III Discussion
Up to now the treatment has been generic for any three-dimensional axionic material in the presence of strong interactions where the axion becomes a fluctuating field. The minimal model of such a dynamical axion insulator is defined by a set of four, anti-commuting elements, that form the basis of the four-dimensional Clifford algebra $Cl_{4}$, and a chiral operator [32, 10]. The phenomenology of this model is equivalent to a correlated Weyl fermion, where the low-energy Hamiltonian is determined by a three-dimensional linear dispersion and a mass with two degrees of freedom, e.g., amplitude and phase. The topology of the ground state at half-filling is characterized by an axion field, which can even become quantized when chiral symmetry is preserved. Within this model, the axion field is intimately related to strong interactions which can endow a dynamical behavior; specific material realizations are given by the corresponding irreducible representations.
For example, in the case of a Weyl semimetal with charge density wave order the form of the Hamiltonian expanded near the degeneracy points is

### v3

@@ Abstract
Introduction—Axionic particles were originally proposed in high-energy physics in order to solve the charge-parity problem as they were associated to the Nambu-Goldstone bosons of a new global axial U(1) symmetry [1, 2]. Today, axions are a leading dark matter candidate and could help explain the matter-antimatter asymmetry present in the universe [3]. While the axion has yet to be discovered in particle physics, its condensed matter counterpart has been theorized to exist in certain three-dimensional materials [4, 5, 6, 7, 8, 9, 10]. A characteristic identifier of the presence of an axion is a signature magnetoelectric response, leading to a modified form of Maxwell’s equations known as axion electrodynamics [11, 12, 13]; these give rise to physical phenomena which can broadly be classified as either transport or density responses. Axion-induced transport includes anomalous Hall currents which are related to the gradient of the axion, and a magnetic conductivity that relates to the axion’s first time derivative. Density responses have to do with the topological magnetoelectric effect where the axion acts as the coupling constant between magnetic field and polarization; and similarly between electric field and magnetization.
In correlated systems, the axion field has an associated collective excitation which couples nonlinearly to the electromagnetic field; this induces a dynamical magnetoelectric effect on top of the mean field response due to the background axion field [14, 15, 16, 17]. A variety of candidate materials have been put forward where axion field effects manifest in the electronic band structure. For example, axion fields naturally appear in correlated Weyl semimetals that undergo lattice translation symmetry breaking due to the emergence of charge density wave (CDW) order. In this case, the mean background axion field is determined by the separation of Weyl points in momentum space and in energy; while its dynamical contribution is associated with collective excitations of the phase of the CDW, i.e., the Nambu-Goldstone bosons [18, 19]. In the presence of uncompensated carrier densities the softening of the axionic collective mode may even lead to spatially inhomogeneous correlated phases [20]. 3D topological insulators (TI) can also support a static, background axion field which, due to time-reversal symmetry, acts as a $\mathbb{Z}_{2}$ topological invariant and takes quantized values $\theta=0$ or $\pi$. When interactions are turned on that spontaneously break time-reversal symmetry, as in the case of antiferromagnetic (AFM) order, $\theta$ is promoted to a dynamical quantity that takes continuous values between 0 and $2\pi$ [5]. Here, the fluctuations of the axion field are induced by the spin-wave excitations of the AFM order parameter [5, 21, 22].
@@
Here we define an all-optical, contactless protocol to identify nonlinear signatures of dynamical axion excitations in a three-dimensional insulating system, termed a dynamical axion insulator (DAI). Our general approach applies to any system hosting axionic excitations. The protocol has two parts: first, a dynamical axion mode is resonantly excited from a background axion field via two-photon Raman processes. By solving the nonlinear coupling between electromagnetic fields and collective axion modes, we show that the axion’s transient evolution depends sensitively on the incident angle, polarization, and frequencies of the two photons. After excitation, a third probe beam detects the resulting axion electrodynamics by monitoring Kerr angle oscillations, serving as a faithful proxy for the axion field’s dynamics [26, 27]. We apply our protocol to a minimal Hamiltonian model that serves as the basis of describing a variety of candidate materials.
Our all-optical, contactless protocol offers advantages over existing methods. Ahn et al. [27] study static, frequency-dependent axions in magnetic TIs requiring broken time-reversal symmetry. In contrast, our method applies to general axionic materials with fluctuating collective modes. While Li et al. [5] and Gooth et al. [23] infer dynamical axions indirectly via static fields or transport, we directly track axion oscillations using time-resolved Kerr rotation—without linearizing the equations of motion—yielding a more general and sensitive probe.
Axion electrodynamics—We start by considering the axion-modified electromagnetic Lagrangian
@@
where $\kappa$ is a materials-dependent parameter, to be given a physical interpretation as “compressibility” later on when we consider the specific case of an axion insulator. We have also included a finite gap for the collective mode $\sim\Omega_{0}$, which reflects the possibility of a resonance at finite frequency, e.g., due to collective mode pinning from disorder.
As we are dealing with electrodynamics, we assume that the speed at which the axion propagates is orders of magnitude slower than the speed of light; as such the axion is effectively dispersionless provided it has a finite gap $\Omega_{0}$. Comparing to the speed of light, $c\gg\nu$ so that all terms relating to first gradients of the dynamical axion field, $\nabla\delta\theta$, can be safely ignored; these describe effects from axion dispersion through the medium.
@@
where we have included a phenomenologically relevant dissipation term $\kappa\gamma\partial_{t}\delta\theta$, with damping rate $\gamma$, which represents the coupling of the order parameter to other degrees of freedom in the bulk. The above equation describes the dynamics of a massive spinless bosonic field, driven by an effective force $\frac{\alpha}{\pi}\bm{E}\cdot\bm{B}$.
@@
Importantly, the axion oscillation amplitude in Eq. (5) depends on: (i) a geometric volume set by polarization vectors $\hat{\bm{\varepsilon}}_{i}$ and momenta $q_{i}/\omega_{i}$, and (ii) a frequency-dependent Lorentzian response. As a second-order nonlinear optical process, excitation occurs via two channels: two-photon absorption (sum or second harmonic generation), Fig.2(a), and stimulated Raman scattering (difference frequency generation), Fig.2(b). The channel dependence arises from the parallelepiped volume formed by $\hat{\varepsilon}_{i}$ and $\bm{q}_{1}/\omega_{1}-\bm{q}_{2}/\omega_{2}$, as shown in Fig.2(c), constraining the response by beam geometry.
The efficacy of this two-photon excitation protocol for a variety of different incident beam angles $\varphi$ and frequencies $\omega_{i}$ in the maximally-crossed polarization geometry, i.e., we keep the polarizations of the excitation beams orthogonal $\hat{\varepsilon}_{1}\cdot\hat{\varepsilon}_{2}=0$, c.f., Fig. 1. In Fig. 3, we show the induced dynamical axion amplitude $|\delta\theta|^{2}$ as a function of incident photon frequencies $\omega_{1},\omega_{2}$ for different incident angles, $\varphi=0,\pi/4$, and $\pi/2$. Fig. 3 clearly shows how varying degrees of freedom of the incident beam can excite the axion field, with the inelastic stimulated Raman channel for collinear beams resulting in the strongest response.
@@
Figure 3: This figure illustrates two distinct axion excitation pathways: two-photon absorption (upper half-plane) and stimulated Raman scattering (lower half-plane), with excitation amplitude $|\delta\theta|^{2}$ shown via color bars. The amplitude is plotted against normalized beam frequencies $\omega_{1}/\Omega_{0}$ and $\omega_{2}/\Omega_{0}$, with $\Omega_{0}=1$, and beam geometries indicated atop each subfigure. (a) For $\varphi=0$ (anti-parallel beams), two-photon absorption dominates. (b) At $\varphi=\pi/2$ (parallel incidence), two-photon absorption is suppressed, and Raman excitation is maximal. (c) For $\varphi=\pi/4$ (orthogonal beams), both channels are present, with Raman excitation strongly enhanced. The grey line at $y=0$ separates the two excitation regimes.
Axion detection—We now move on to demonstrate how the axion collective mode may be detected. The key idea is to use a linearly polarized detection beam of known frequency at normal incidence, which when reflected off the surface of the material will posses a time-dependent Kerr effect induced by the presence of the dynamical axion  [26].
@@
The dispersion relation of electromagnetic fields in the material is obtained by solving the equations of motion of the electromagnetic field $\epsilon\partial^{2}_{t}{\bm{A}}-\frac{1}{\mu}\nabla^{2}\bm{A}=\partial_{t}\bm{J}$, where $\epsilon$ is the relative permittivity, $\mu$ is the relative permeability, and $\bm{J}$ is the current given in Eq. (2). Without loss of generality we assume a non-magnetic, spatially isotropic medium such that the permeability and permittivity tensors are simplified to $\mu^{-1}=\mathds{1}$ and $\epsilon=\epsilon\mathds{1}$, where $\epsilon$ is a constant characterizing the material. Furthermore, we consider propagation along the z-axis and solve the equations of motion in the plane orthogonal to the direction of propagation. The refractive index given by (see S.I.) $n_{\pm}=\left(\frac{\epsilon\pm\epsilon^{\prime}}{1\pm\mu^{\prime}}\right)^{1/2}$, with corresponding dispersion $k=\omega n_{\pm}$, where $\epsilon^{\prime}=\omega^{-1}{\partial_{z}\theta}$ is related to the electric conductivity and $\mu^{\prime}=k^{-1}{\partial_{t}\theta}$ is related to the magnetic conductivity.
The reflectance and transmittance amplitudes can be obtained via the continuity of electric and total magnetic fields at the interface, given as $\bm{E}^{I}+\bm{E}^{R}=\bm{E}^{T}$, and $\bm{B}^{I}+\bm{B}^{R}=\bm{B}^{T}+\frac{\alpha}{\pi}\delta\theta\bm{E}^{T}$ (see S.I.). Without loss of generality, here we only include the dynamical axion $\delta\theta$ contributions to the Fresnel equations, and set the background axion at the boundary to zero. Any such contributions, as in $\mathds{Z}_{2}$ topological insulators where the axion takes quantized values, can be incorporated in the equilibrium position of the dynamical axion mode $\delta\theta$. Using the matching conditions, the coefficients of reflection are computed as
@@
where $R_{\pm}=\frac{A_{\pm}^{R}}{A_{\pm}^{I}}$ is the ratio between the reflected and initial amplitudes of the fields. The reflection coefficients are, in general, different depending on the refractive index of left- and right-circularly polarized light, and the dynamical axion angle.
By linearizing Eq. (6) for small $\delta\theta$ one obtains the following expression for the polar Kerr rotation angle:
$\Theta_{K}\approx\frac{\alpha}{\pi}\frac{(-2+n_{+}^{2}+n_{-}^{2})}{(n_{+}^{2}-1)(n_{-}^{2}-1)}\delta\theta.$ (7)
The final result of this work is contained within the above equation, which shows the modulation of the reflected beam’s polar Kerr rotation angle and its dependence on the collective fluctuations of the axion field, c.f., Eq. (5). This serves as our smoking gun optical signature to unambiguously deduce the presence of a dynamical axion field in the material.
@@
Figure 4: Transient Kerr rotation reveals signatures of excited axion modes. Using a pump profile $\bm{E}\cdot\bm{B}\propto e^{(t/\sigma)^{2}}\cos(\Omega_{D}t)$ tuned to two-photon resonance ($\Omega_{D}=\Omega_{0}$), we drive out-of-phase axion oscillations $\delta\theta(t)$ with relative phase $\pi/2$ set by damping $\gamma$. These oscillations are reflected in the time-dependent Kerr angle $\Theta_{K}(t)$, which closely follows the axion dynamics. Amplitudes are illustrative and not to scale.
Material application—Up to now the treatment has been generic for any three-dimensional axionic material in the presence of strong interactions where the axion becomes a fluctuating field. The minimal model of such a dynamical axion insulator is defined by a set of four, anti-commuting elements, that form the basis of the four-dimensional Clifford algebra $Cl_{4}$, and a chiral operator [28, 10]. The phenomenology of this model is equivalent to a correlated Weyl fermion, where the low-energy Hamiltonian is determined by a three-dimensional linear dispersion and a mass with two degrees of freedom, e.g., amplitude and phase. The topology of the ground state at half-filling is characterized by an axion field, which can even become quantized when chiral symmetry is preserved. Within this model, the axion field is intimately related to strong interactions which can endow a dynamical behavior; specific material realizations are given by the corresponding irreducible representations.
For example, a Weyl semimetal with charge density wave order may host dynamical axion modes. In an uncorrelated Weyl semimetal with no CDW the nodes are gapless and the chirality of the Weyl nodes is given by the topological charge associated to each band crossing, which is identical to the Berry flux $C=+1$ (C=1C=-1) for the right- (left-) handed Weyl particles [29] for a spinful system. However, in the presence of sufficiently strong interactions, a transition to a CDW ordered phase with complex order parameter $\Delta$ can occur.
Due to the chiral anomaly, it is known that the phase of the complex CDW can be identified with the axion response [6] of the material. The total CDW phase $\theta$ is given by $\theta(\bm{r},t)=\bm{Q}\cdot\bm{r}+Q_{0}t+\delta\theta(\bm{r},t)\,$, where the first two terms are dubbed the background axion field $\theta_{0}=\bm{Q}\cdot\bm{r}+Q_{0}t$ and are determined by the band structure of the material: the separation in momentum-space $\bm{Q}$, and energy $Q_{0}$, between the Weyl nodes. The former is non-zero only in the case of time-reversal symmetry breaking, while the latter is non-zero for inversion-symmetry breaking systems. The final term $\delta\theta(\bm{r},t)$ is the dynamical axion mode, which can be understood as the collective mode associated to the sliding phase mode of the CDW. In this case, the parameter $\kappa$ in Eq. (3) is physically interpreted as the chiral compressibility.